# Enhancing Wide-Angle Image Using Narrow-Angle View of the Same Scene

### v2

@@ Abstract
A common dilemma while photographing a scene is whether to capture it at a wider angle, allowing more of the scene to be covered but in less detail or to click in a narrow angle that captures better details but leaves out portions of the scene. We propose a novel method in this paper that infuses wider shots with finer quality details that is usually associated with an image captured by the primary lens by capturing the same scene using both narrow and wide field of view (FoV) lenses. We do so by training a Generative Adversarial Network (GAN)-based model to learn to extract the visual quality parameters from a narrow-angle shot and to transfer these to the corresponding wide-angle image of the scene using residual connections and an attention-based fusion module. We have mentioned in details the proposed technique to isolate the visual essence of an image and to transfer it into another image. We have also elaborately discussed our implementation details and have presented the results of evaluation over several benchmark datasets and comparisons with contemporary advancements in the field.
@@ 1 Introduction
Modern smartphones and mirrorless cameras increasingly incorporate ultra-wide lenses [1] to capture expansive scenes in a single frame. While these lenses deliver a striking field of view, they often struggle to preserve the fine-grained detail achieved by conventional “primary” lenses, as discussed in the paper [2]. Users face a perennial trade-off: opting for a wider perspective at the expense of sharpness or retaining detail by narrowing the view. In this work, we seek to resolve this dilemma by synthesizing wide-angle images whose perceptual quality matches that of narrow-angle shots, without sacrificing the original field of view.
Our approach hinges on jointly leveraging both narrow- and wide-angle captures of the same scene. The narrow-angle image supplies high-resolution texture and detail, while the wide-angle image carries context beyond the primary lens’s range. We introduce an end-to-end pipeline that incrementally transfers visual fidelity from the finest (narrowest) view through to the widest view in a multi-lens stack. First, we partition both the wide image and the corresponding narrow-angle image into spatially aligned patches inspired by the paper on Vision Transformer [3]. Each patch pair is fed into a generator, which fuses content and detail via cross-view attention and further super-resolves the wide patch using pixel-shuffle layers—progressively enhancing resolution and texture across the full frame.
Central to our method is a novel GAN [4] architecture: a generator tailored to fuse cross-view content and detail features through attention mechanisms and a discriminator designed to enforce both global realism and seamless patch blending. We further extend the classic Gram matrix representation to encode not only texture but also color, brightness, contrast, and sharpness metrics, ensuring rich visual fidelity. Applying this procedure iteratively across all available zoom levels—from telephoto to ultra-wide—we end up with a single image that combines the sweeping perspective of the widest lens with the crisp detail of the primary lens. Our key contributions are
@@ 2 Related Works
Another stream of relevant research comes from neural style transfer. Initially developed for artistic applications [9], these models demonstrated that image attributes could be transferred from one image to another. Subsequent work improved efficiency using feed-forward networks [10] and adaptive normalization techniques [11], broadening their use cases. However, transferring quality between narrow and wide FOV images introduces unique challenges not fully addressed by style transfer methods, as the wide FOV often contains additional contextual elements that require accurate integration.
@@
Recent advancements have introduced lightweight and efficient architectures tailored for real-time applications. LatticeNet [12] employs a lattice block structure inspired by filter banks, enabling adaptive feature fusion while maintaining low computational complexity. This design facilitates the reconstruction of high-quality images without the overhead associated with deeper networks. Building upon the success of attention mechanisms, the Efficient Long-Range Attention Network (ELAN) [13] integrates shift convolutions with group-wise multiscale self-attention modules. This combination captures both local and long-range dependencies, enhancing the model’s ability to reconstruct fine details across varying scales. The ELAN architecture achieves a balance between performance and efficiency, making it suitable for applications that require real-time processing.
Further exploring attention-based models, DLGSANet [14] introduces a hybrid approach combining dynamic local and sparse global self-attention mechanisms. The Multi-Head Dynamic Local Self-Attention (MHDLSA) module effectively captures local features, while the Sparse Global Self-Attention (SparseGSA) module focuses on the most relevant global information. This dual attention strategy enhances the model’s capacity to reconstruct high-resolution images with reduced computational demands. In parallel, the Deep Multilevel Feature Fusion Network (DMFFN) [15] addresses common challenges in super-resolution tasks, such as blurred edges and redundant structures. Using nested two-level residual blocks with asymmetric structures, DMFFN efficiently fuses multilevel features, leading to improved edge preservation and faster convergence during training. These contemporary models demonstrate significant progress in balancing performance and efficiency. Their architectures are designed to maintain structural consistency and fine details, addressing the limitations observed in earlier methods.
Perceptual metrics such as LPIPS [16] have become widely adopted for evaluating visual similarity in a way that aligns closely with human perception. Beyond perceptual loss, incorporating visual cues that guide perception—namely color, brightness, contrast, and sharpness—is equally important. We select these attributes based on extensive literature: [17] identifies them as key appearance properties in image quality assessment; [18] highlights the role of color in quality estimation; [19] emphasizes the critical balance of brightness and contrast in enhancement; and [20] combines contrast and sharpness to assess enhancement reliability.
By combining insights from generative models, reference-based super-resolution, attention mechanisms, and perceptual consistency techniques, we target high-quality, structurally accurate image synthesis suited for immersive visual applications.
@@ 3 Method
The core philosophy of our solution is to use a cross-view attention mechanism to fuse content and visual quality features from a given pair of wide and narrow-angle images. The images are initially partitioned into fixed-size patches. We pick a pair of wide and narrow patches from the pool and pass them across the network architecture, as described below.
@@ 3.1 Image Partitioning
Both narrow and wide FoV images are partitioned into $128\times 128$-sized patches to ensure portions of images sharing visual similarity are used to exchange content and visual quality data. Once partitioned, patches from wide and narrow FoV images that share visual properties are mapped using perceptual loss [21] and stored in the pool, ready to be pushed into the generator network. A cosine similarity on the feature maps generated by the mid layers of a VGG-19 [22] backbone is used to determine the perceptual distance between patches.
@@ 3.2 Network Architecture
The architecture of the generator network is built upon a stack of ResNet-like [23] residual blocks that take as inputs a pair of wide and narrow FoV patches from the pool. The patches are encoded using separate encoder components to capture structural (content) and visual features from wide and narrow FoV patches respectively. Each block then employs a cross-view attention mechanism to fuse the content and visual properties from wide and narrow feature maps respectively. The pipeline consists of the following major components,
@@ 3.2.2 Visual Feature Encoder
These encoders are specialized in extracting the visual parameters from a patch and are used to encode the narrow FoV patches. We use the Gram matrix approach to represent the visual features of a patch and include the following additional parameters in the Gram matrix representation: color, brightness, contrast, and sharpness. The feature map is constructed as follows,
  • The original feature map $F_{l}\in\mathbb{R}^{C_{l}\times H_{l}\times W_{l}}$ is generated using a VGG-19 network.
  • Color information for each channel over spatial dimensions, also referred to as global average pooling [24], is computed as:
    $[F_{l_{\text{color}}}]_{c}$ $=\frac{1}{H_{l}W_{l}}\sum_{h=1}^{H_{l}}\sum_{w=1}^{W_{l}}F_{l}(c,h,w),$
    for $\quad\text{for }c=1,\dots,C_{l}$
  • Global mean pixel intensity across all channels and spatial dimensions indicates the perceived luminance (brightness) of the patch. Computed as:
    $F_{l_{\text{brightness}}}=\frac{1}{C_{l}H_{l}W_{l}}\sum_{c=1}^{C_{l}}\sum_{h=1}^{H_{l}}\sum_{w=1}^{W_{l}}F_{l}(c,h,w)$
  • We represent the overall contrast of the patch using the standard deviation:
    $F_{l_{\text{contrast}}}=\sqrt{\frac{1}{C_{l}H_{l}W_{l}}\sum_{c=1}^{C_{l}}\sum_{h=1}^{H_{l}}\sum_{w=1}^{W_{l}}\left(F_{l}(c,h,w)-\mu_{l}\right)^{2}}$
    where $\mu_{l}=F_{l_{\text{brightness}}}$ is the global mean defined above.
  • Sharpness is accounted for using the Laplacian operated, as mentioned in this paper [25]:
    $F_{l_{\text{sharpness}}}(c,h,w)=(F_{l}(c,\cdot,\cdot)*\nabla^{2})(h,w)$
    where * denotes the convolution and $\nabla^{2}$ is the Laplacian kernel.
  • Finally, we augment the original feature map obtained from VGG-19 as follows:
    $\hat{F}_{l}=[F_{l},F_{l_{\text{color}}},F_{l_{\text{brightness}}},F_{l_{\text{contrast}}},F_{l_{\text{sharpness}}}]$
    where scalar or vector features are broadcast to match the spatial dimensions of $F_{l}$.
Once the enhanced feature map has been computed, we perform a dot product between the enhanced feature map and its transpose to attain the Gram matrix representation $E_{n}$, carrying the visual essence of the narrow FoV patch.
$E_{n}=\hat{F}_{l}\hat{F}_{l}^{\top}$
@@ 3.2.3 Cross-Attention Residual Blocks
Once the patches are encoded into the required latent space, we employ an attention-based fusion mechanism that fetches the most relevant visual features from the narrow patch and fuses those with the content features generated from the wide patch.
@@
  • Encoded narrow FoV patch $E_{n}$, carrying the visual properties, provides both the key and value matrices
  • The attention scores, which gather semantically aligned, superior quality visual features present in the encoded narrow FoV patch, are computed using the scaled dot product equation introduced in the attention paper [26],
    $Attention(E_{n},E_{w})=softmax(\frac{E_{w}E_{n}^{\top}}{\sqrt{d_{k}}})E_{n}$
  • The gathered attention scores are then incorporated into the wide FoV feature map using a residual addition, producing the feature map representation of the enhanced wide FoV patch, $E$
    $E=E_{w}+Attention(E_{n},E_{w})$
@@ 3.2.4 Upsampling Layers
The output of the final residual block is passed through a pixel shuffle [27] layer with a scale factor equal to 2 to super-resolve the patch, denoted by $E^{*}$.
@@ 3.3 Loss Calculation
Generator loss is calculated patch-wise, i.e., each of the generated patches is compared against both the original wide and narrow FoV patches to quantify the content and visual deviations from the ground truth. Then the patches are assembled using seamless cloning, to eliminate visible seams during blending, and the reconstructed image is passed through the discriminator which computes a seam consistency loss over the entire image. The losses are described as follows,
@@ 3.3.1 Generator Loss
    $L^{(l)}_{content}(E^{o}_{w},E^{g}_{w})=\frac{1}{n}\sum_{i}(E^{o}_{w}-E^{g}_{w})^{2}$
    Here $E^{o}_{w}$ and $E^{g}_{w}$ are the latent space representations of the original and generated patches.
@@
    Visual Quality Loss: This loss is calculated as the difference of the Gram matrix representation of the generated patch and the narrow FoV patch using the following equation,
    $L^{(l)}_{visual}=\sum_{l}w_{l}||E^{o}_{n}-E^{g}_{n}||^{2}$
    Here $E^{g}_{n}$ and $E^{o}_{n}$ are the gram matrix representations of the patches from the generated and the original narrow FoV patches, respectively.
@@ 3.3.2 Discriminator Loss
    Seam Consistency Loss This loss is aimed at eliminating the seams between the patches in the synthetic image by minimizing the perceptual difference at the boundaries of the adjacent patches. For adjacent patches $P_{i}$ and $P_{j}$, we take neighboring regions $R_{i}$ and $R_{j}$ which are $w$ pixels wide and compute the loss as,
    $L^{(l)}_{seam}=\frac{1}{w}\sum||\phi_{l}(R_{j})-\phi_{l}(R_{i})||^{2}$
@@
    Perceptual Loss against Ground Truth To ensure the essence of the narrow FoV has been plausibly transferred to the synthetic image, the output is perceptually compared to the original wide FoV ground truth. We use the same VGG19 activation maps to find the perceptual difference $L^{(l)}_{perceptual}$ between the generated wide FoV image and the original image from the dataset.
@@ 4.1 Datasets
We have used Landscapes Dataset [28] (LHQ 1024×1024) to train the proposed model. With 90,000 high-quality images of 1024 x 1024 resolution, it includes a diverse range of urban and natural environments. These high-quality landscape images with rich textures and details challenge the method’s ability to handle complex natural scenes. We have randomly sampled 20000 + 500 images (discussed in the Training Details section) and fed into the training loop in the ascending order of pixel variance [29], to ensure “easier” images are encountered first.
@@
[H] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] 2x Zoom 1x Zoom 1x Output [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] 1x Output 0.5x Zoom 0.5x Output
@@ 4.2 Simulating Narrow and Wide FoVs
4.3 Generator
The generator takes in a pair of patches, each of dimension $128\times 128\times 3$. The wide FoV patch is passed through a CNN encoder with a final batchNorm layer to produce the content feature map. The narrow FoV patch initially undergoes feature extraction through the $Conv2_{2},Conv3_{2},Conv4_{2}$ layers of a VGG-19 backbone. The feature map produced by VGG-19 is enhanced by the visual properties computed from the original patch. We use 10 attention-based residual blocks, each with two $3\times 3$ convolution units, a ReLU, followed by the cross-view attention score calculation unit, outputs of which are added to the wide FoV patch. The details-infused wide FoV patch is then passed through a series of convolution layers interleaved by pixel shuffle layers to physically upscale the patch to a dimension of $256\times 256\times 3$.
4.4 Discriminator
We use a ResNet18 [31] based discriminator network that takes as input a $1024\times 1024\times 3$ image formed by stitching the enhanced patches generated by the generator using Poisson Blending [32]. The image undergoes a series of $6$ residual blocks followed by a layer of adaptive pooling and a fully connected layer outputting a single scalar denoting the realness of the input image.
4.5 Enhanced image reconstruction pipeline
Figure 5 highlights the steps undertaken to reconstruct the enhanced.

### v3

@@ Abstract
A common dilemma in photography is the trade-off between capturing a wider field of view with less detail and a narrower field of view with finer detail. To address this, we propose a method that enhances wide-angle images by leveraging co-captured narrow-angle views of the same scene. Our approach introduces a narrow field-of-view (FoV) image-informed enhancement strategy that overcomes the limitations of blind super-resolution, which often struggles to recover high-frequency details without auxiliary guidance. We train a GAN-based model to extract fine-grained visual features from the narrow-view image and fuse them into the corresponding wide-view image through residual connections and an attention-based feature integration module. This fusion is guided by perceptual cues representing the visual essence—such as color, contrast, and sharpness—learned from the narrow FoV. We describe our architectural design and implementation in detail and demonstrate the effectiveness of our method through extensive experiments and comparisons with recent state-of-the-art techniques.
@@ 1 Introduction
Modern smartphones and mirrorless cameras increasingly incorporate ultra-wide lenses [1] to capture expansive scenes in a single frame. While these lenses deliver a striking field of view, they often struggle to preserve the fine-grained detail achieved by conventional “primary” lenses. As discussed in the paper [2], this difficulty is faced due to various technical issues, such as barrel distortion and spreading the view across a greater number of pixels, resulting in a lower perceived resolution. Users face a perennial trade-off: opting for a wider perspective at the expense of sharpness or retaining detail by narrowing the view. In this work, we seek to mitigate this dilemma by synthesizing wide-angle images whose perceptual quality matches that of narrow-angle shots, without sacrificing the original field of view.
While traditional super-resolution (SR) algorithms can increase the spatial resolution of an input image, they must rely on statistical priors and learned patterns to hallucinate high-frequency details, often leading to inconsistent results and visual artifacts. Our approach, on the other hand, hinges on jointly leveraging both narrow and wide-angle captures of the same scene. The narrow-angle image supplies high-resolution texture and detail, hence supplanting the blind dependence on learned distributions, while the wide-angle image carries context beyond the primary lens’s range. The enhancement objective, in this paper, is formulated as a resolution boosting task, informed by fine detail cues from the narrow FoV image, allowing us to draw upon proven strategies and validation procedures from SR literature. To that end, we introduce an end-to-end pipeline that utilizes cross-view attention scores to pinpoint the relevant features in the narrow FoV image embedding that can be used to guide the upscaling of the wide FoV counterpart. Inspired by the paper on Vision Transformer [3], we adopt a patch-based approach to ensure the quality transfer occurs between regions that share visual similarity.
Central to our method is a proposed GAN [4] architecture: a generator attempting to fuse cross-view content and detail features through attention mechanisms and a discriminator that aims at enforcing both global realism and seamless patch blending. We further extend the classic Gram matrix representation to encode not only texture but also color, brightness, contrast, and sharpness metrics, ensuring rich visual fidelity. Applying this procedure iteratively across all available zoom levels, from telephoto to ultra-wide, our method aims to synthesize the broad perspective of the widest lens while improving detail quality using information from the primary lens. Our key contributions are,
@@ 2 Related Works
Another stream of relevant research comes from neural style transfer. Initially developed for artistic applications [9], these models demonstrated that image attributes could be transferred from one image to another. Subsequent work improved efficiency using feed-forward networks [10] and adaptive normalization techniques [11], broadening their use cases. However, transferring quality between narrow and wide FOV images introduces unique challenges not addressed by style transfer methods, as the wide FOV often contains additional contextual elements that require accurate integration.
@@
Recent work includes LatticeNet [12], which uses lattice-style blocks for adaptive feature fusion with low computation, enabling high-quality reconstruction without deep networks. However, its reliance on localized operations limits its capacity to capture long-range dependencies and global context, making it less effective in scenes with large structural variation or complex spatial arrangements. ELAN [13] combines shift convolutions with group-wise multiscale self-attention to capture both local and long-range dependencies, achieving a balance of accuracy and efficiency for real-time tasks. Nevertheless, the use of fixed attention windows introduces boundary discontinuities and restricts the model’s ability to fully capture cross-region interactions, particularly in non-uniform or geometrically diverse image regions.
Further exploring attention-based models, DLGSANet [14] employs a hybrid of Multi-Head Dynamic Local Self-Attention (MHDLSA) for local features and Sparse Global Self-Attention (SparseGSA) for key global cues, improving high-resolution reconstruction with lower compute. While this hybrid structure improves reconstruction quality, it models local features implicitly and may introduce irrelevant global tokens, leading to suboptimal detail preservation in fine-grained textures or highly cluttered scenes. Similarly, DMFFN [15] employs nested two-level asymmetric residual blocks to fuse multilevel features, enhancing edge preservation and accelerating convergence. However, its fixed fusion design limits adaptability across spatial scales, and deeper residual nesting can introduce training instability. These architectures, while advancing performance-efficiency trade-offs, may still face challenges in maintaining structural consistency under complex geometric variations.
To address these limitations, we propose a cross-view attention-based framework that explicitly models global context, aligns geometric structure across views, and adaptively fuses multi-scale features, leading to more robust reconstruction under challenging structural and visual conditions. By combining insights from generative models, reference-based super-resolution, attention mechanisms, and perceptual consistency techniques, our approach targets high-quality, structurally accurate image synthesis suited for immersive visual applications.
@@ 3 Method
We define the task of perceptually enriching wide-angle imagery as a guided resolution enhancement problem. The model receives a pair of co-captured images: a wide field-of-view (FoV) image that offers spatial context, and a narrow FoV image that retains fine-grained visual fidelity. The output is a physically upscaled version of the wide FoV image, enriched with perceptual cues derived from the narrow FoV capture. While the model performs a spatial upscaling operation on the wide FoV input, the true goal is not just to increase pixel count, but to enrich the image perceptually. The resolution increase provides a larger canvas, enabling the injection of meaningful details such as sharp textures, balanced contrast, and natural colors. A detailed, step-by-step walk-through of the methodology is provided below.
@@ 3.1 Image Partitioning
We start by dividing both the narrow and wide FoV images into fixed-sized patches to ensure portions of images sharing visual similarity are used to exchange content and visual quality data. Once partitioned, patches from wide and narrow FoV images that share visual properties are mapped using perceptual loss [16] and stored in the pool, ready to be pushed into the generator network. A cosine similarity on the feature maps generated by the mid layers of a VGG-19 [17] backbone is used to determine the perceptual distance between patches. During inference, as further explained in the experimentation section, the same partitioning ratio is maintained.
@@ 3.2 Network Architecture
The architecture of the generator network is built upon a stack of ResNet-like [18] residual blocks that take as inputs a pair of wide and narrow FoV patches from the pool. The patches are encoded using separate encoder components to capture structural (content) and visual features from wide and narrow FoV patches, respectively. Each block then employs a cross-view attention mechanism to fuse the content and visual properties from wide and narrow feature maps, respectively. The pipeline consists of the following major components and has been visually illustrated in Fig 1.
@@ 3.2.2 Visual Feature Encoder
Narrow FoV patches are passed through encoders that are specially crafted to capture the visual essence of an image. We intend to do so by initially computing the Gram matrix representation of the patch and then augmenting it with visual cues computed from the patch. The augmented Gram matrix is then termed as $E_{n}$. As mentioned earlier, we take the color, brightness, contrast, and sharpness parameters of a patch as the representative of its perceptual indicators. Detailed mathematical formulations are included in the supplementary material(Section 1).
@@ 3.2.3 Cross-Attention Residual Blocks
Once the patches are encoded into the required latent space, we employ an attention-based fusion mechanism that fetches the most relevant visual features from the narrow patch and fuses those with the content features generated from the wide patch. This step facilitates the propagation of high-frequency details into the resultant image.
@@
  • Encoded narrow FoV patch $E_{n}$, carrying the visual properties, provides both the key and value matrices. $d_{n}$ is the dimension of the matrix $E_{n}$
  • The attention scores, which gather semantically aligned, superior quality visual features present in the encoded narrow FoV patch, are computed in equation (1), using the scaled dot product equation introduced in the attention paper [19],
    $Attention(E_{n},E_{w})=softmax(\frac{E_{w}E_{n}^{\top}}{\sqrt{d_{n}}})E_{n}$ (1)
  • The gathered attention scores are then incorporated into the wide FoV feature map using a residual addition in equation (2), producing the feature map representation of the enhanced wide FoV patch, $E$
    $E=E_{w}+Attention(E_{n},E_{w})$ (2)
@@ 3.2.4 Upsampling Layers
The output of the final residual block is passed through a pixel shuffle [20] layer, spatially upscaling the patch to accommodate the incorporated perceptual cues, denoted by $E^{*}$. Generator architecture has been illustrated in Fig 3
@@ 3.3 Loss Calculation
Generator loss is calculated patch-wise, i.e., each of the generated patches is compared against both the original wide and narrow FoV patches to quantify the content and visual deviations from the ground truth. Then the patches are assembled using seamless cloning to eliminate visible seams during blending, and the reconstructed image is passed through the discriminator, which computes a seam consistency loss over the entire image. The losses are described as follows,
@@ 3.3.1 Generator Loss
    $L^{(l)}_{content}(E^{o}_{w},E^{g}_{w})=\frac{1}{n}\sum_{l}(E^{o}_{w}-E^{g}_{w})^{2}$ (3)
    Here $E^{o}_{w}$ and $E^{g}_{w}$ are the latent space representations of the original and generated patches, and $n$ is the dimension.
@@
    Visual Quality Loss: This loss is calculated as the weighted sum of differences of the Gram matrix representations of the generated patch and the narrow FoV patch, from mid layers of the VGG-19 backbone, using the following equation,
[…]
    Here $E^{g}_{n}$ and $E^{o}_{n}$ are the Gram matrix representations of the patches from the generated and the original narrow FoV patches, respectively. $w_{l}$ is the weight assigned to the layer $l$.
@@ 3.3.2 Discriminator Loss
    Seam Consistency Loss This loss is aimed at eliminating the seams between the patches in the synthetic image by minimizing the perceptual difference at the boundaries of the adjacent patches. For adjacent patches $P_{i}$ and $P_{j}$, we take neighboring regions $R_{i}$ and $R_{j}$ which are $b$ pixels wide and compute the loss as,
[…]
@@
    Perceptual Loss against Ground Truth To ensure the essence of the narrow FoV has been plausibly incorporated into the synthetic image, the output is perceptually compared to the original wide FoV ground truth. We use the same VGG19 activation maps to find the perceptual difference $L^{(l)}_{perceptual}$ between the generated wide FoV image and the original image from the dataset.
@@ 4.1 Datasets
We have used Landscapes Dataset [21] (LHQ 1024×1024) to train the proposed model. With 90,000 high-quality images of 1024 x 1024 resolution, it includes a diverse range of urban and natural environments. These high-quality landscape images with rich textures and details provide a rigorous training ground for the model to learn how to handle complex natural scenes. We have randomly sampled 20000 + 500 images (discussed in the Training Details section) and fed them into the training loop in the ascending order of pixel variance [22], to ensure “easier” images are encountered first. Furthermore, to accommodate the rectangular aspect ratios of ultra-wide images encountered in real-world scenarios, we have fine-tuned the model on the DIV2K dataset [23], which consists of 800 images at 2K (1920 × 1080; 16:9) aspect ratios, thereby facilitating the model’s ability to generalize effectively across varying aspect ratios.
@@
[H] Narrow Input Wide Input Output [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] 2x Zoom 1x Zoom 1x Output [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] 1x Output 0.5x Zoom 0.5x Output
@@ 4.2 Simulating Narrow and Wide FoVs
4.3 Image Reconstruction Pipeline
During training, each image is passed through a reconstruction pipeline. The cycle begins with loading a pair of images, wide and narrow FoV, from the dataset, applying the preprocessing mentioned earlier, and dividing the processed images into patches. The patches, from the pool, are then fed into the generator, which produces their enhanced counterparts. The enhanced patches are stitched back into a single image, the super-resolved image, and forwarded to the discriminator for loss computation. This pipeline has been illustrated in Fig 5. Implementation details about the generator and discriminator components are provided in the supplementary document (Section 2).